OpenStudy (anonymous):

alex has $20, and ellen has $12. alex is saving $3 per day, and ellen is saveing $5 per day. after how many days will alex and ellen have the same amount of money

OpenStudy (anonymous):

@texaschic101 please help

OpenStudy (charlotte123):

Alex - 20 ^ Saves 3 dollars per Day Ellen - 12 ^ Saves 5 dollars per Day Alex ----> 20 + 3x Ellen -----> 12 + 5x

OpenStudy (anonymous):

?

OpenStudy (zzr0ck3r):

set those two equations equal to each other

OpenStudy (charlotte123):

^

OpenStudy (charlotte123):

20 + 3x = 12 + 5x ----> 3x + 20 = 5x + 12 -----> (3x + 20) + (-20 - 5x) = (5x + 12) + (-20 - 5x) Thus - x = 4

OpenStudy (texaschic101):

20 + 3x = 12 + 5x 20 - 12 = 5x - 3x 8 = 2x 8/2 = x 4 = x I completely agree with charlotte :)

OpenStudy (texaschic101):

wait...you got a -x ??

OpenStudy (charlotte123):

@texaschic101 *-*

OpenStudy (charlotte123):

@texaschic101 No no LOL A Dash - xD

OpenStudy (texaschic101):

ok....that makes more sense...lol

OpenStudy (charlotte123):

My bad LOL

OpenStudy (anonymous):

thanks

OpenStudy (texaschic101):

sure thing :) thats what we are here for

OpenStudy (charlotte123):

No problem! @NaomiBell1997